OpenStudy (anonymous):

Integrate: (dy)/(dx)= (xy)/(x^2-1)

OpenStudy (anonymous):

here is the question as written from Diff EQ, verify that x^2+cy^2=1 where c is an arbitrary nonzero constant, is one parameter family of implicit solutions to \[\frac{ dy}{ dx }=\frac{ xy }{ x^2-1 }\] So, I found that the original equation is \[y=\sqrt{\frac{ 1-x^2 }{ c }}\]

OpenStudy (anonymous):

So how do I begin to integrate from here, It has been a while.

OpenStudy (anonymous):

Would y be a arbitrary number and be able to be brought out of the integral?

OpenStudy (anonymous):

When you differentiate \(y\) did you get \(dy/dx\)?

OpenStudy (anonymous):

The original problem you do not have to differentiate. Its basically proving it.

OpenStudy (anonymous):

It has already been done.

OpenStudy (anonymous):

So what exactly is the question asking for? \( \int y dx\)?

OpenStudy (anonymous):

Here is where I am at with the integral \[\int\limits_{}^{}\frac{ xy }{ x^2-1 } \frac{ dy }{ dx}\] \[Let u=x^2-1 \] \[du=2x dx\]

OpenStudy (anonymous):

Thats why I don't know if I can take the y out of the integral and solve. I have not used this calculus in quite some time.

OpenStudy (anonymous):

If you do a \(u\) sub like that, then you'd have to find \(y\) in terms of \(u\).

OpenStudy (anonymous):

If I wanted to solve for \(y\), I would separate the function.

OpenStudy (anonymous):

the equation I mean. It's a separable equation.

OpenStudy (anonymous):

I need to show the original problem \[x^2+cy^2=1\]

OpenStudy (anonymous):

Okay, you can either use implicit differentiation, or you can just separate the equation and integrate.

OpenStudy (anonymous):

With algebra \[y=\sqrt{\frac{ 1-x^2 }{ c }}\]

OpenStudy (anonymous):

Okay sorry, but you have totally confused me now. You mention integration, you have a derivative, but you only want to use algebra...? 1) What is given exactly? 2) What do we need to prove/show exactly?

OpenStudy (anonymous):

problem written as above. Section is Solutions and Initial Values Problems.

OpenStudy (anonymous):

How about integration by parts, where \(u = y\)?

OpenStudy (anonymous):

and \(dv = x/(x^2-1)\)?

OpenStudy (anonymous):

@nestor18 did you try it?

OpenStudy (anonymous):

working it had to refresh on ibp

OpenStudy (tkhunny):

Why are we trying to integrate in one stroke a separable differential equation? \(\int \dfrac{1}{y}\;dy = \int\dfrac{x}{x^{2}-1}\;dx\)

OpenStudy (anonymous):

I am not sure, but when I mentioned separation, @nestor18 didn't want to do it.

OpenStudy (tkhunny):

@wio Fair enough. I see that on re-reading. @nestor18 Simplify your life. Solve as a separable differential equation. Seriously.

OpenStudy (anonymous):

I am so rusty with this stuff, I forget all rules for integration. This makes more sense now.

OpenStudy (tkhunny):

\(\ln|y| = \dfrac{1}{2}\ln\left(|x^{2}-1|\right) + C\) Show us the rest.